OpenStudy (anonymous):

Refer to triangle DEF. If m Ang D = 30°, m Ang E = 60°, and ME = 4, what is DM in simplest form? 4 8 4 sqrt(3) 12

OpenStudy (squirrels):

Do you have an attachment for this question?

OpenStudy (anonymous):

i thought i posted it woops sry http://www.learningbygrace.org/geom/images/geom_24.1.JPG

OpenStudy (squirrels):

kk one moment (:

OpenStudy (anonymous):

thank you i will most defiantly give you a medal

OpenStudy (squirrels):

D. 12 (:

OpenStudy (squirrels):

tan 60 = FM/ME = FM/4 => FM = 4* (root 3) FM/DM= tan 30 => DM = 4(root 3) * (root 3) = 4*3 = 12

OpenStudy (squirrels):

And thats how ya do it

OpenStudy (anonymous):

thank you very much sir ^_^